1) Gardener A can mow an entire lawn in 3 hours. Gardener B can mow the same lawn in 5 hours. How long will it take them to mow the lawn if they work together?
2)To paint a house alone it would take Mario 3 days, Danny 4 days and Lito 6 days. How many days should it take them to paint 3 houses if they all work together?

Answer by nerdybill(7384)

1) Gardener A can mow an entire lawn in 3 hours. Gardener B can mow the same lawn in 5 hours. How long will it take them to mow the lawn if they work together?
Let x = time (hours) it takes working together
then
x(1/3 + 1/5) = 1
multiplying both sides by 15:
x(5 + 3) = 15
x(8) = 15
x = 15/8 hours
or
x = 1.875 hours
or
x = 1 hour and 52.5 minutes
.
2)To paint a house alone it would take Mario 3 days, Danny 4 days and Lito 6 days. How many days should it take them to paint 3 houses if they all work together?
Let x = time (days) it takes working together
then
x(1/3 + 1/4 + 1/6) = 3
multiplying both sides by 12:
x(4 + 3 + 2) = 36
x(9) = 36
x = 4 days